OpenStudy (newwar):

Solve for p. 3+|1+p|=12 p = 8 or p=−10 p = 14 or p=−16 p = 10 or p=−10 p = 8 or p=−8

OpenStudy (mathmale):

Start by isolating |1+p|. Do this by subtracting 3 from both sides of your equation. Show your work, please.

OpenStudy (newwar):

3+1=4 so D

563blackghost (563blackghost):

Not quite. We would isolate \(\large{|1+p|}\) so we would subtract 3 from both sides in order to get it to itself so.... \(\huge{(3-3)|1+p|=12-3}\) What would that equal?

OpenStudy (newwar):

(3-3=0) (1+p=?) (12-3=9)

OpenStudy (mathmale):

\[\huge{(3-3)|1+p|=12-3}\]

OpenStudy (mathmale):

is not correct. Both of you, please work on this to find and correct the error.

OpenStudy (mathmale):

This post is newwar's; please help newwar solve the problem himself or herself.

OpenStudy (mathmale):

Again, subtract 3 from both sides of the original equation.

OpenStudy (brainzonly):

I'm following mathmale around giving him medals :D

OpenStudy (mathmale):

newwar? 3+|1+p|=12 -3 = -3 -------------- Please complete this work.

563blackghost (563blackghost):

I seemed to have written the equation wrong but I seem to have gotten the right answer. Sorry about that.

OpenStudy (newwar):

i am sorry i dont know what to do first

OpenStudy (mathmale):

Given 3+|1+p|=12 -3 = -3 -------------- please begin by subtracting 3 from 3. Type out your result.

OpenStudy (mathmale):

3 - 3 ----- ??

OpenStudy (newwar):

that is 0

OpenStudy (mathmale):

True, so then you have |1+p| on the left side. What have you on the right side?

OpenStudy (newwar):

subtract 3

OpenStudy (mathmale):

Yes, subtract 3 from 12. Your result?

OpenStudy (newwar):

-9